OpenStudy (lilai3):

Write an equation in slope-intercept form of the line satisfying the given conditions: is perpendicular to 4y=3x-8 and passes through (0,-6) @xxyyxyyx

OpenStudy (anonymous):

the slope of a line that is perpendicular to another line is the negative reciprocal of that line

OpenStudy (anonymous):

first write the given equation in slope intercept form, so simpifying you will get y=3/4x-2

OpenStudy (anonymous):

an example of that is if slope = 2, then the slope of a perpendicular line is -1/2

OpenStudy (anonymous):

since you know from point intercept formula: y=mx+c, m = slope of the line, comparing the above simplified equation you can see m = 3/4

OpenStudy (anonymous):

so the slope of the above line is 3/4

OpenStudy (anonymous):

but the new line that we need to write the eqation of is perpendicular to the above line

OpenStudy (anonymous):

there is a formula for the slope of perpenicular lines. m2= -1/m1

OpenStudy (anonymous):

where m2 is the slope of the perpendicular line, and m1 is the slope of the given line

OpenStudy (anonymous):

so inputting the values we get. m2= -1/(3/4)

OpenStudy (anonymous):

so m2= -4/3

OpenStudy (lilai3):

so every perpendicular line is the opposite and the reciprocal? like for example... -1/3 → would become, if perpendicular, → 3/1 or 3.... is that what you mean?

OpenStudy (anonymous):

Now you have the slope of the perpendicular line. you can write the euation of the line like in the previous question

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

since this time you only have one point given in the question (0,-6) input them and simplify

OpenStudy (anonymous):

and rearrange and write the answer in y=mx+c form

OpenStudy (anonymous):

so you will get y-(-6)=-4/3(x-0) y+6=-4/3x y=-4/3x-6

OpenStudy (anonymous):

yup parallel lines will have same slope. m2=m1

OpenStudy (anonymous):

and perpendicualar lines will have m2=-1/m1

OpenStudy (anonymous):

negative reciprocal of the given line

OpenStudy (lilai3):

wait.. shouldn't the perpendicular lines have m2=-1/m?

OpenStudy (lilai3):

be***, not have***

OpenStudy (anonymous):

yes

OpenStudy (lilai3):

oh.... sorry, you could continue...

OpenStudy (lilai3):

negative reciprocal of the given line...

OpenStudy (anonymous):

yes m2=-1/m1 means m2 = negative reciprocal of m1

OpenStudy (anonymous):

where m1 is the given line

OpenStudy (lilai3):

okay...although this is sorta confusing.....it's hard to explain where it is confusing, it just is... oh phooey

OpenStudy (anonymous):

k let me know if you have other stuffs

OpenStudy (lilai3):

anyway, i get it now, after a few tweeks and fr-leeks. thanks so much! you're the best! medal rewarded x2!!!! thank you!

OpenStudy (anonymous):

your welcome. what does the medal do btw.

OpenStudy (lilai3):

hmm...i'm not really sure. it's sort of a "thank you" kind of thing. it also helps you raise your level (: LOL sorry for the late response.